Question Help

The human resource department at a certain company wants to conduct a survey regarding worker benefits. The department has an alphabetical list of all

51695169

employees at the company and wants to conduct a systematic sample of size

6060.

​(a) What is

k​?

kequals=86 86

​(b) Determine the individuals who will be administered the survey. Randomly select a number from 1 to k. Suppose that we randomly select

1414.

Starting with the first individual​ selected, the individuals in the survey will
Answer #1

Ans:

a)

k=5169/60=86

b)

j= 20.

So,the units in the sample will be 14, 100,186,................,5088.

(last unit=14+86*(60-1)=5088)